OpenStudy (anonymous):

A meal at a restaurant costs $10.00 and the tax, yet to be added is 7.5%, plus you wish to leave a tip of 15% of the untaxed bill. The total amount that you would pay (tax and tip included), is $12.36. [a] True [b] False

OpenStudy (anonymous):

When i added all of them together i got 12.36 so i think it's true

OpenStudy (mathstudent55):

What is 7.5% of $10?

OpenStudy (mathstudent55):

75% of $10 = 7.5/100 * 10 =

OpenStudy (anonymous):

who are you asking

OpenStudy (mathstudent55):

Whoever is interested in learning how to answer this question.

OpenStudy (anonymous):

well 7.5% of $10 is 0.75 and then you have 10.75

OpenStudy (mathstudent55):

Before we actually do calculations, let's see what we need to do. We have a $10 bill (no tax & no tip) We need to add 7.5% tax and 15% tip. We need to find 7.5% of $10 (the tax) and we need to find 15% of $10 (the tip) and then we add $10, the tax and the tip.

OpenStudy (mathstudent55):

Great. So far, you already found that the the tax is $0.75, and you added the tax, so the meal + tax is $10.75 Now we need to find the tip. The tip is 15% of $10. You can do this 2 ways. 1. Just find 15% of $10 15% of 10 = 15/100 * 10 = 2. Notice that 15% is exactly 2 times the 7.5% tax, so the tip is twice the tax. 2 * 0.75 = Both ways will give the same amount.

OpenStudy (mathstudent55):

The tip ends up being $1.50 $10.75 + $1.50 = $12.25, not $12.36

OpenStudy (mathstudent55):

Correct. Now you add $10.75 + $1.5 What do you get?

OpenStudy (anonymous):

12.25 wow you are so smart

OpenStudy (anonymous):

^_^

OpenStudy (mathstudent55):

Thanks. As you can see, the answer is: [b] False